OpenStudy (anonymous):

Identify the center and the radius of the circle with equation x2 + (y - 5)2= 9 center (0, 5) and radius = 3 center (0, 5) and radius = 81 center (0, -5) and radius = 81 center (0, -5) and radius = 3

OpenStudy (anonymous):

use equation of the circle again :)

OpenStudy (anonymous):

its blah+blah=r^2

OpenStudy (anonymous):

if your r^2=9 then r must be?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

x2 + (y - 5)2= 9 is already in the equation

OpenStudy (anonymous):

A or D :D

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 9\implies 3^2}} \\ \quad \\ \quad \\ x^2 + (y - 5)2= 9\implies (x-0)^2 + (y - 5)2= 3^2\) what do you think?

OpenStudy (anonymous):

+5 according to equation

OpenStudy (jdoe0001):

woops, darn. typo \(\bf {\color{blue}{ 9\implies 3^2}} \\ \quad \\ \quad \\ x^2 + (y - 5)2= 9\implies (x-0)^2 + (y - 5)^2= 3^2\)

OpenStudy (jdoe0001):

\(\bf x^2 + (y - 5)^2= 9\implies (x-0)^2 + (y - 5)^2= 3^2\) any ideas?

OpenStudy (anonymous):

It's A I think.

OpenStudy (jdoe0001):

\(\bf (x-{\color{red}{ 0}})^2 + (y - {\color{blue}{ 5}})^2= 3^2\Leftarrow center=(0,5)\qquad radius=3\) yeap

OpenStudy (anonymous):

ah thank you!! uvu

OpenStudy (jdoe0001):

yw